What is the value of A+B if the A is the fist term and the B is the common difference of the sequence whose 4th term is 3 and 20th term is 35?
Answer by Maths68(1474)

an = a1 + (n – 1)d
n=4
an=3
3=a1+(4-1)d
3=a1+3d........(1)
an = a1 + (n – 1)d
n=20
an=35
35=a1+(20-1)d
35=a1+19d........(2)
Subtract (1) from (2)
35=a1+19d
-3=-a1-3d
----------------
32=16d
32/16=d
2=d
d=2
Put the value of d in (1)
3=a1+3d........(1)
3=a1+3(2)
3=a1+6
3-6=a1
-3=a1
First term = A =-3
Common Difference = B = 2
A+B=-3+2=-1